OpenStudy (anonymous):

what are the three next terms 3,8,13,18

OpenStudy (anonymous):

Do you want to learn how to setup an equation to these points or do it the old fashion way of seeing a pattern and doing it in your head??

OpenStudy (anonymous):

i want to learn how to setup an equation

OpenStudy (anonymous):

n 1 2 3 4 3 8 13 18 5 5 5 Since the difference is a constant number in the first line of taking the differences, we know the equation is just constant times n instead of it being squared. The constant here is 5. So we now know 5n. If we take the first term 5(1) + b = 3, we find that b is -2. So our equation for this series is 5n-2

OpenStudy (anonymous):

Use this equation to solve for the next three terms

OpenStudy (anonymous):

Did you follow my explanation?

OpenStudy (anonymous):

i tried but i dont get it

OpenStudy (anonymous):

First step you find the difference

OpenStudy (anonymous):

if the difference is constant right away then you are done. And the equation takes the form of (constant)n + b = y

OpenStudy (anonymous):

if it is constant after taking the difference twice then it will be (constant)n^2

OpenStudy (anonymous):

This is a Ap with Common differnce 5 ie T2 - T1 = 8-3 = 5

OpenStudy (anonymous):

So...next term is to add 5 to the last term ie 18 +5 = 23

OpenStudy (anonymous):

ok thankyou :)